# A bi-fidelity method for the uncertain Vlasov-Poisson system near quasineutrality in an asymptotic-preserving particle-in-cell framework

### v2

@@ 1. Introduction
The novelty of this work is two-fold: (i) The deterministic solver for the VPME system, namely the AP-PIC method developed and numerically implemented is new, inspired by yet different from the earlier work [47] where the macroscopic equations are studied. Our designed numerical scheme in the PIC framework satisfies the asymptotic-preserving property and can capture the quasineutral limit on the discrete level. (ii) This is the first work on studying the VPME system with multi-dimensional random uncertainties, by employing the multi-fidelity approach. We point out that in each problem setting, the choice of the low-fidelity model can be very different and is not unique, which directly affects the accuracy and efficiency of the bi-fidelity approximation. In the second part of our work, the novel part of solving UQ problems is to choose the macroscopic Euler system as the low-fidelity solver, while using the AP-PIC method developed in (i) as a robust high-fidelity solver.
@@ 3.2. Derivation of the reformulated Poisson equation
We note that although the original Poisson equation (3.13) and the reformulated Poisson equation (3.14) are equivalent, they are not equally suited in the quasi-neutral limit. In fact, the quasi-neutral potential equation (3.18) automatically serves as the formal limit $\varepsilon\to 0$ of the reformulated Poisson equation (3.14), but not of the original Poisson equation (3.13) which gives $n_{i}=e^{\phi}$ as $\varepsilon\to 0$. Therefore, in order to design asymptotic-preserving schemes, it is necessary to base on the reformulated system (3.17)–(3.18) for numerical discretization. We will introduce our AP scheme in the PIC framework in the next section.
@@ A.1. Derivation of isothermal compressible Euler system
This closure assumption for the ion moment system can be found in [39, Section 2]. Based on this closed system, one considers the Euler-Poisson-Boltzmann (EPB) model [47] by neglecting the ion pressure tensor, where the cold-ion approximation (i.e., $T_{i}\to 0$, so $\tau\ll 1$) is adopted:

### v3

@@ 1. Introduction
The novelty and main contribution of our work is two-fold: (i) In this work, inspired by [17], where the reformulation idea is introduced, we design a PIC method for the VPME system that includes a semi-implicit particle update and an implicit reformulated Poisson equation. Our proposed scheme is asymptotic-preserving and can capture the correct quasi-neutral limit, while allowing the time step and mesh size to be much larger than the Debye length. Moreover, our AP-PIC scheme, built based on reformulation, decouples the field-particle iteration, where the particle trajectories and macroscopic moments are computed only once per time step, significantly improving computational efficiency compared with a fully implicit PIC scheme [12]. We also mention that compared with the work by Degond et al. [48] that considers the macroscopic system, our model involves the Vlasov equation on the kinetic level, thus their implicit scheme can not be directly adapted here. (ii) To our best knowledge, this is the first work to study the UQ problem for the VPME system. We explore different multi-fidelity approaches using appropriately chosen low-fidelity models and evaluate their accuracy and efficiency. For the high-fidelity solver, the AP-PIC method developed in (i) is adopted, which further underscores the importance of achieving high efficiency for our deterministic solver.
@@ 3.2. Derivation of the reformulated Poisson equation
We note that although the original Poisson equation (3.13) and the reformulated Poisson equation (3.14) are equivalent, they are not equally suited in the quasi-neutral limit.
@@ A.1. Derivation of isothermal compressible Euler system
This closure assumption for the ion moment system can be found in [40, Section 2]. Based on this closed system, one considers the Euler-Poisson-Boltzmann (EPB) model [48] by neglecting the ion pressure tensor, where the cold-ion approximation (i.e., $T_{i}\to 0$, so $\tau\ll 1$) is adopted: